OpenStudy (anonymous):

I'm trying to make sure my answer is correct. Subtract: 3/x - 7/(x-2)

OpenStudy (anonymous):

\[\frac{3}{x}-\frac{7}{x-2}=\frac{3(x-2)-x\times7}{x(x-2)}\] is a start

OpenStudy (anonymous):

I got that

OpenStudy (anonymous):

rest is distributive law and combining like terms in numerator. don't look to cancel!

OpenStudy (anonymous):

I have -2(2x+3)/x(x-2)

OpenStudy (anonymous):

i get \[\frac{3x-6-7x}{x(x-2)}=\frac{-4x-6}{x(x-2)}\]

OpenStudy (anonymous):

yes you are correct!

OpenStudy (anonymous):

Yayyy! Thank you! I just like checking with other people to see if my answer is correct and if they got the same thing.

OpenStudy (anonymous):

good work!